OpenStudy (nthenic_oftime):

any help would be great :) ... If the force acting on an object stays the same, then the acceleration of the object is inversely proportional to its mass. If an object with a mass of 9 kilograms accelerates at a rate of 6 meters per second per second by a force, find the rate of acceleration of an object with a mass of 2 kilograms that is pulled by the same force. A. 3 meters per second per second B. 27 meters per second per second C. 18 meters per second per second D. 24 meters per second per second

OpenStudy (baru):

hi

OpenStudy (nthenic_oftime):

hello @baru

OpenStudy (welshfella):

Use Newtons second law of motion Force = mass * acceleration

OpenStudy (nthenic_oftime):

D?

OpenStudy (baru):

@nthenic_oftime have you started with physics...or is this a problem from math class?

OpenStudy (nthenic_oftime):

@baru i know some physics bt this is from math

OpenStudy (welshfella):

F = 9*6 = 54 N for the mass of 2 kg:- 54 = 2 a

OpenStudy (nthenic_oftime):

thank you @welshfella

OpenStudy (welshfella):

sole for a

OpenStudy (welshfella):

yw

OpenStudy (nthenic_oftime):

wait wouldnt i have to do 2*6 because its asking for object with mass of 2kg at same force

OpenStudy (welshfella):

the force is 54 N 6 is the acceleration for the 9 kg mass

OpenStudy (welshfella):

acceleration for 2 kg mass = 54/2

OpenStudy (baru):

@welshfella i think this is an exercise in ratios...not physics..he/she is not very familiar with force

OpenStudy (nthenic_oftime):

ooh okay get it we found the force first and then found rate of acceleration of the 2kg mass gotchya

OpenStudy (nthenic_oftime):

i dont understand why this is on a math test when to me it seems like physics

OpenStudy (welshfella):

well physics is often 'applied math'.

OpenStudy (baru):

you are supposed to understand how "inverse proportionality" works

OpenStudy (welshfella):

yea when a is inversely proportional to m we can write a = k / m where k is some constant in the above case the constant is the force (54)

OpenStudy (welshfella):

you can find the value of k by plugging in known values of a and m ( 6 and 9) 6 = k/9 so k = 6*9 = 54 know we have established the formula for a in terms of m a = 54/m so knowing the value of a or m we can find the other variable

OpenStudy (welshfella):

so when m = 2 a = 54/ 2

OpenStudy (nthenic_oftime):

okay @welshfella thats what i was seeing in my mind or what made sense i didnt know it had a name thank you for that info too

OpenStudy (nthenic_oftime):

lol

OpenStudy (welshfella):

with inverse variation as one variable increses the other decreases and vice versa